OpenStudy (anonymous):

A thrift shop sells costumes for d dollars. Today they are at a rate of $7 each. Maryellen purchased 3 dicounted costumes and a wig that costs $12. What is the original costs of the costumes if Maryellen spends a total of $51?

OpenStudy (agent47):

if she spent 51 in total, and $12 was spent on a wig, then she spent: $51-$12=$39 on 3 discounted costumes

OpenStudy (agent47):

$39/3 = $13 on each costume.

OpenStudy (anonymous):

I need the work cant understand it with out the work pease

OpenStudy (agent47):

So total that was spent = $51 $12 of that $51 was spent on a wig, so the rest must have been spent on costumes

OpenStudy (agent47):

Thus, she spent: $51-$12=$39 on the costumes

OpenStudy (anonymous):

what about the discount

OpenStudy (agent47):

I'm not done yet.. She spent $39 on 3 costumes, so she spent $39/3=$13 on each.

OpenStudy (anonymous):

so d=39$?

OpenStudy (anonymous):

what grade r u in

OpenStudy (agent47):

This question is wrong.... discounted price cannot be more than the original price..

OpenStudy (agent47):

aa820 that question is irrelevant

OpenStudy (anonymous):

the discount was at a rate of 7$

OpenStudy (agent47):

ohhhhhh.. so you wrote the question wrong..

OpenStudy (anonymous):

oops sorry lol

OpenStudy (agent47):

You said the costumes were $7, no, the discount is $7

OpenStudy (anonymous):

yea i meat the dicount wa $7

OpenStudy (agent47):

So if she bought them for $13, with a $7 discount, they should've cost $7 more, or $13+$7=$20

OpenStudy (agent47):

basically if she bought something for $A at $B discount, then the original price was $(A+B)

OpenStudy (anonymous):

let me rewrite the whole ? again i think iwrote it wrong

OpenStudy (anonymous):

A thrift shop sells costumes for d dollars. Today they are discounted at a rate of $7 each. Maryellen purchased 3 dicounted costumes and a wig that costs $12. What is the original costs of the costumes if Maryellen spends a total of $51?

OpenStudy (anonymous):

that should be right

OpenStudy (anonymous):

Okie, original price is 13 + 7 = $20